OpenStudy (anonymous):

Find the volume of a cylinder with diameter 22 in and height 15 in. Use 3.14 for pi and round to the nearest hundredth

OpenStudy (anonymous):

V(cylinder) = 2πr.h or V(cylinder) = πd.h V(cylinder) = 3.14 x 22 x 15 V(cylinder) = 1036.20 cubic inches

OpenStudy (anonymous):

@Silent_Sorrow, I cannot remain silent and say with great sorrow that the volume of a cylinder is the area of the base (not the perimeter) times the height: V=Pi*R^2*H

OpenStudy (anonymous):

Oh crap, too right you are @CarlosGP. My apologies to all. V=Pi*R^2*H V=3.14 x (22/2)^2 x 15 V=3.14 x 11^2 x 15 V=3.14 x 121 x 15 V=5699.1 cubic inches